This is not acually for this topic, I need help on adding, subtracting, multiplying and dividing positive and negative numbers. How do I do this?
Example; -23- 73=???

Answer by richwmiller(17219)

-23-73=-96
Think of it as adding two regular numbers but they both have a minus sign.
I am sure you can do 73-23=50 and -23+73=50
also you can think of it as I owe John 73 cents and I owe Mary 23 cents
James owes me 15 cents
-73-23+15=-93+15=-78
What you owe is negative (going out)and what they owe you is positive. (coming in)